OpenStudy (anonymous):

The sum of a number and twice its reciprocal is 27/5. find the number

OpenStudy (anonymous):

x + 2(1/x) = 27/5

OpenStudy (anonymous):

okay but then what?

OpenStudy (anonymous):

and is it 1/x because thats the reciprocal?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so then x + 2/x + 27/5

OpenStudy (anonymous):

= 27/5

OpenStudy (anonymous):

multiply everything by x to get rid of the fraction. x^2 + 2 = 27/5x

OpenStudy (anonymous):

5x^2 -27x + 10 =0

OpenStudy (anonymous):

then factor to solve it i guess or use the quadratic formula. I am not 100% sure of this, double check or see if the answer u get is correct

OpenStudy (anonymous):

idk how to factor that. i cant do it

OpenStudy (anonymous):

idk i guess thats the only way to do it

OpenStudy (anonymous):

(5z-2)(z-5)=0